if a number is decreased by its reciprocal, the result is 11 over 30. find the number.
Answer by venugopalramana(3286)

if a number ..X...SAY....is decreased by its reciprocal....1/X... the result is 11 over 30. find the number.
X-1/X=11/30........MULTIPLY WITH 30X THROUGHOUT
30X^2-30=11X
30X^2-11X-30=0
30X^2-36X+25X-30=0
6X(5X-6)+5(5X-6)=0
(5X-6)(6X+5)=0
5X-6=0....THAT IS 5X=6....X=6/5...
OR
6X+5=0...THAT IS 6X=-5....X=-5/6..